luhivqqcherry:

help me . 1. 4x + y - 5h = 10y + u, for x 2. 12x + 8 - 15 = -2 (3x -82) 3. 45x - 720 + 15x = 60

Kagura:

1. X= u/4 + 5h/4 + 9y/4 2. X= 19/2 3. X= 13 @luhivqqcherry

luhivqqcherry:

@kagura wrote:
1. X= u/4 + 5h/4 + 9y/4 2. X= 19/2 3. X= 13 @luhivqqcherry
THANK YOU SO MUCH < 3

Kagura:

np